OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y =1/3 x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (ajprincess):

The gradient of the given line is 1/3. Let the gradient of the required line be m. Then \[m*\frac{1}{3}=-1\] from this m can be found. \[gradient~m=\frac{y-y_1}{x-x_1}\] \[m(x-x_1)=y-y_1\] \[y=m(x-x_1)+y_1\] \[y=mx-mx_1+y_1\] Here \(x_1=4~and~y_1=-2\) Does that help? @manueld167